# Worldsheet patching, 1-form symmetries, and “Landau-star” phase transitions

### v1

@@ I Introduction
It is a remarkable fact that many topological phase transitions, which lack any local order parameter, can be described using Landau-Ginsburg theory anyway. The Landau theory is formulated in terms of a “fictitious” order parameter, which does not exist as a local observable of the original theory. Canonical examples include Higgs transitions in discrete gauge theories, and in three dimensions also their confinement transitions [1, 2]. The simplest case is the 3D gauge-Higgs model with $\mathbb{Z}_{2}$ gauge field and $\mathbb{Z}_{2}$ matter, where both the Higgs and confinement transitions are “$*{}^{*}$” transitions [3, 4, 5, 6, 7, 8]. These are not Ising transitions, but they can be largely understood in terms of the latter. In the deep infra-red, the $*{}^{*}$ fixed point, or more generally a “$*{}^{*}$” fixed point, is related to the standard Landau fixed point by orbifolding: i.e. by a gauging procedure that eliminates non-gauge-invariant operators, including the Landau order parameter, from the spectrum. (We will review ways of thinking about $*{}^{*}$ transitions below.)
A crucial point about the relationship between the gauge theory and the Landau theory is that in general it is emergent, rather than microscopic. In other words, the relation is meaningful only after coarse-graining beyond some characteristic lengthscale $\ell$. For a Higgs transition, $\ell$ separates a regime at shorter scales, where gauge fluctuations are nontrivial, from a regime at larger scales where (heuristically) the gauge field can be treated as flat — i.e. locally equivalent to pure gauge. Once we have an effective model where nontrivial gauge fluctuations have been eliminated, the relation to a Landau-like theory is direct: we can locally choose a gauge where the gauge field is trivial, so that the interactions of the Higgs fields resemble those in a Landau theory [1, 2]. The emergence of the $*{}^{*}$ description is also tied to the emergence of a “higher” symmetry [9, 10, 11, 12, 13], associated with emergent string operators, as discussed below.
@@
In the case where $\ell$ is suppressed all the way to zero, the “duality” between the transition in the gauge theory and the Landau transition holds at the microscopic level [1]. (This may be extended to small $\ell$ within perturbation theory [1, 2].) However, the scale $\ell$ may be arbitrarily large; for example we can approach a phase transition where $\ell$ diverges (and where the $*{}^{*}$ description breaks down). In general the $*{}^{*}$ description and the associated higher symmetry emerge only after a nontrivial renormalization group flow.
@@
First, to develop a geometrical understanding of how the $*{}^{*}$ description — with its “fictitious” order parameter — emerges as we go from the microscopic scale to scales $\gg\ell$. In a more formal language, this is also the emergence [12, 14, 15, 16, 17, 18] of a one-form symmetry [9, 10, 11] — a symmetry group generated by string operators. (Here, these string operators are associated with anyons [19, 13, 12].) We start from the geometrical representation of the partition function of a lattice gauge theory as an ensemble of membranes [20, 21, 22, 23, 14] and use a “patching” procedure to construct the emergent observables. The general idea was outlined in Ref. [14] and is developed here. The procedure is not limited to lattice gauge theory, and could be applied to many other statistical mechanics ensembles formulated in terms of strings or membranes (we will discuss some examples).
@@
The patching approach also sheds light on topological phase transitions that are not $*{}^{*}$-like and which are far less understood, including the self-dual point where the two $*{}^{*}$ lines in Fig. 1 meet [25, 26, 14], which is itself a scale-invariant critical point [14, 27, 28]. We discuss the fate of the string operators at this critical point. Interestingly, although they become nonlocal, there is a sense in which this nonlocality is weak.
@@
Figure 1: Phase diagram of the $\mathbb{Z}_{2}$ gauge–Higgs model in 3D, showing the confinement transition line (red), the Higgs transition line (black), the self-dual multicritical point (MCP) where they meet, and a segment of first order coexistence extending from the MCP (double line).
@@
In the 3D $\mathbb{Z}_{2}$ gauge-Higgs model, the Higgs and confinement lines (shown in Fig. 1) are completely equivalent, by the exact duality property of the model [1]. In our convention, the transition line we discuss is the confinement transition. As can be seen in Fig. 1 we can cross this line at various values of the Higgs coupling, which is related to $y$ in the figure. As we move along the confinement phase transition line by increasing $y$, the typical size $\ell$ of the “holes” in the membrane increases, diverging at $y_{\text{MCP}}$, the self-dual critical point where Higgs and confinement lines meet [14, 20]. We study the confinement ($*{}^{*}$) transition at various values of $y$, up to and including $y_{\text{MCP}}$.
@@ II.1 Definition of the model
There are two dual ways to formulate the membrane mapping. The equivalence of these dual representations is one way to see the equivalence of the Higgs and confinement phase transition lines in (Fig. 1). In a quantum language, these dual formulations correspond to formulations of the path integral either in the “electric” or the “magnetic” basis, as discussed below. We choose the “electric” representation, in which the membrane ensemble is
@@
Eqs. 1 and 3 can be thought of as 3D classical partition functions in their own right. However, if we choose to think of $Z$ as an imaginary time path integral for a 2+1D quantum $\mathbb{Z}_{2}$ gauge theory, then the membrane ensemble (3) is the formulation of this path integral in the basis of $\mathbb{Z}_{2}$ electric flux lines which live in the 2D spatial plane and which terminate on electric charges. The membranes we are discussing are worldsurfaces of these electric flux lines (see e.g. Fig. 6 in Ref. [14]). The boundaries of the membranes, which make up the “holes”, are worldlines of the electric charges. Duality is the existence of a completely equivalent picture in the magnetic basis, for $\mathbb{Z}_{2}$ magnetic flux lines and charges. [^3] The dual (i.e. magnetic) membrane ensemble is a straightforward rewriting of Eq. 2: a plaquette of the dual cubic lattice is occupied if ${\tau\sigma\tau=-1}$ for the corresponding link of the original lattice.
@@ II.3 Finite and infinite holes
Heuristically, when $\ell(y)$ is finite (${y<y_{\text{MCP}}}$), we might imagine that holes disappear under RG. That is, after coarse-graining to scales ${\gg\ell(y)}$ we effectively recover a theory of closed membranes, and the system flows to the same $*{}^{*}$ fixed point that governs the phase transition on the boundary ${y=0}$. Below we will show that this can be made concrete (giving a constructive mapping to Ising) also for ${y>0}$.
@@ III.2 Membrane-patching algorithm
Having defined the patches we can define the repaired configuration $\overline{\mathcal{M}}$. If we call the set of patching plaquettes ${\cal P}$, the new set ¯$\overline{\mathcal{M}}={\cal M}\cup{\cal P}-{\cal M}\cap{\cal P}$ defines strictly closed surfaces and allows us to construct line operators as described in Sec. III.1.
@@ III.3 Ising observables
As discussed around Eq. 4, an ensemble of closed membranes can be split into sectors. Our patching algorithm works in all sectors, but for some observables we will restrict to the first sector (${\sigma_{x}=\sigma_{y}=\sigma_{z}=+}$). In this sector the dual Ising spin $S_{r}$ has periodic boundary conditions and so is well-defined up to a global sign flip, so that after patching the correlation function $\left\langle S_{r}S_{r^{\prime}}\right\rangle$ may be straightforwardly defined. When we write expressions such as $\left\langle S_{r}S_{r^{\prime}}\right\rangle$, we refer to an average conditioned on the ${\sigma_{x}=\sigma_{y}=\sigma_{z}=+}$ sector.
(More formally, this may be expressed as
@@
where $P$ is any path connecting $r$ to $r^{\prime}$, and $\chi_{+}(\mathcal{M})$ is an indicator function that takes the value $\chi_{+}=1$ if $\mathcal{M}$ is a patchable configuration in the +++++++ + + sector, and the value $\chi_{+}=0$ otherwise.)
@@
III.4 Diagnosing the deconfined phase
Since in Sec. IV we will focus on critical scaling near the confinement transition line, we emphasize here that the patching procedure gives a way to characterize the deconfined phase more generally (i.e. even far from the confinement transition line). The deconfined phase is the part of the phase diagram where both ($i$) and ($ii$) hold:
($i$) The patching procedure succeeds: i.e. the probability of a configuration $\mathcal{M}$ being patchable tends to 1 at large system size. Equivalently, the holes are non-percolating.
($ii$) The dual Ising order parameter is disordered, as manifested, for example, in the vanishing of $\left\langle M^{2}\right\rangle$ at large $L$. Equivalently, the expectation value of the string operator ${\lim_{L\to\infty}\langle\widetilde{V}_{P}\rangle}$, for an open path of length $R$, tends to zero at large $R$.
@@ IV.1 Dual Ising correlators
A striking demonstration of $*{}^{*}$ universality is given by the power-law scaling of the dual Ising correlator as a function of separation $z$:
@@
The correlator is shown in Panel 1 of Fig. 5, as a function of distance $z$ in a system of fixed size $L=64$. Data for the values ${y=0,0.1,0.2}$, which lie on the confinement line and are expected to be in the $*{}^{*}$ universality class, are consistent with the Ising power law, where $\eta_{\mathrm{Ising}}=0.036298(2)$ [31], which is shown with a dotted line. [^7] The power law applies for $1\ll r\ll L$. It is also possible to perform a scaling collapse for $r/L$ of order 1 (not shown). (Since the Ising anomalous dimension $\eta_{\text{Ising}}$ is small, this line is close to $1/z$.) The distinct curves are also consistent with each other within error bars: although in general the nonuniversal normalization of $C(z)$ will depend on $y$, this dependence appears to be very weak. The Panel also shows the correlator computed at the MCP, where its RG interpretation is different and will be discussed in Sec. V.2.
@@ IV.2 Other sectors
At the $*{}^{*}$ critical point, and in the limit $L\rightarrow\infty$, these numbers are universal and related by duality to ratios of Ising partition functions with different boundary conditions,
@@
In our Monte Carlo dynamics, the configuration can make a transition between different sectors by passing through a non-patchable configuration with a percolating hole. In principle this allows equilibration over all sectors. In practice this is possible if $y$ is not too small and $L$ is not too large. When $y<y_{\text{MCP}}$ such transitions are exponentially suppressed in the limit $L\rightarrow\infty$ simply because $P_{\text{perc}}(L)$ is exponentially suppressed. For this reason we have not tried to accurately compute the sector probabilities for the $*{}^{*}$ critical point. This could be done, using the present gauge theory, by parallel tempering, exchanging configurations at different values of $y$ [32].
@@ V.1 Case where $\ell$ is finite
An alternative way to quantify locality in the dual language would be to reexpress the partition function in terms of $S_{r}$.111111 Separating $Z$ into topological sectors (see previous sections) we write e.g. $Z_{+++}=\sum_{\mathcal{M}\in(+++)}e^{-\mathcal{H}[\mathcal{M}]}$ (for the +++++++ + + sector; other sectors are similar) as $Z_{+++}=\frac{1}{2}\sum_{\{S_{r}\}}\sum_{\mathcal{M}\in(+++)}e^{-\mathcal{H}[\mathcal{M}]}\chi(S,\mathcal{M})$, where $\chi(S,\mathcal{M})=1$ if $S$ is one of the two spin configurations consistent with the patched configuration $\overline{\mathcal{M}}$ and zero otherwise. Then the effective spin Hamiltonian is given by $e^{-\mathcal{H}^{\prime}[S]}=\sum_{\mathcal{M}}e^{-\mathcal{H}[\mathcal{M}]}\chi(S,\mathcal{M})$. Because of patching the effective interactions would be nonlocal, but could be approximated by interactions of finite range $\mathcal{O}(\ell)$ at the cost of incurring only a small error in the free energy density.

### v2

@@ I Introduction
It is a remarkable fact that many topological phase transitions, which lack any local order parameter, can be described using Landau-Ginsburg theory anyway. The Landau theory is formulated in terms of a “fictitious” order parameter, which does not exist as a local observable of the original theory. Canonical examples include Higgs transitions in discrete gauge theories, and in three dimensions also their confinement transitions [1, 2]. The simplest case is the 3D gauge-Higgs model with $\mathbb{Z}_{2}$ gauge field and $\mathbb{Z}_{2}$ matter, where both the Higgs and confinement transitions are “Ising” transitions [3, 4, 5, 6, 7, 8]. These are not Ising transitions, but they can be largely understood in terms of the latter. In the deep infra-red, the Ising fixed point, or more generally a “Landau” fixed point, is related to the standard Landau fixed point by orbifolding: i.e. by a gauging procedure that eliminates non-gauge-invariant operators, including the Landau order parameter, from the spectrum. (We will review ways of thinking about Landau transitions below.)
A crucial point about the relationship between the gauge theory and the Landau theory is that in general it is emergent, rather than microscopic. In other words, the relation is meaningful only after coarse-graining beyond some characteristic lengthscale $\ell$. For a Higgs transition, $\ell$ separates a regime at shorter scales, where gauge fluctuations are nontrivial, from a regime at larger scales where (heuristically) the gauge field can be treated as flat — i.e. locally equivalent to pure gauge. Once we have an effective model where nontrivial gauge fluctuations have been eliminated, the relation to a Landau-like theory is direct: we can locally choose a gauge where the gauge field is trivial, so that the interactions of the Higgs fields resemble those in a Landau theory [1, 2]. The emergence of the Landau description is also tied to the emergence of a “higher” symmetry [9, 10, 11, 12, 13], associated with emergent string operators, as discussed below.
@@
In the case where $\ell$ is suppressed all the way to zero, the “duality” between the transition in the gauge theory and the Landau transition holds at the microscopic level [1]. (This may be extended to small $\ell$ within perturbation theory [1, 2].) However, the scale $\ell$ may be arbitrarily large; for example we can approach a phase transition where $\ell$ diverges (and where the Landau description breaks down). In general the Landau description and the associated higher symmetry emerge only after a nontrivial renormalization group flow.
@@
First, to develop a geometrical understanding of how the Landau description — with its “fictitious” order parameter — emerges as we go from the microscopic scale to scales $\gg\ell$. In a more formal language, this is also the emergence [12, 14, 15, 16, 17, 18] of a one-form symmetry [9, 10, 11] — a symmetry group generated by string operators. (Here, these string operators are associated with anyons [19, 13, 12].) We start from the geometrical representation of the partition function of a lattice gauge theory as an ensemble of membranes [20, 21, 22, 23, 14] and use a “patching” procedure to construct the emergent observables. The general idea was outlined in Ref. [14] and is developed here. The procedure is not limited to lattice gauge theory, and could be applied to many other statistical mechanics ensembles formulated in terms of strings or membranes (we will discuss some examples).
@@
The patching approach also sheds light on topological phase transitions that are not Landau-like and which are far less understood, including the self-dual point where the two Ising lines in Fig. 1 meet [25, 26, 14], which is itself a scale-invariant multicritical point [14, 27, 28]. We discuss the fate of the string operators at this multicritical point (MCP). Interestingly, although they become nonlocal, there is a sense in which this nonlocality is weak. We discuss why this MCP is challenging to analyze using continuum field theory (despite the existence of a continuum Chern-Simons formulation for $\mathbb{Z}_{2}$ topological order).
@@
Figure 1: Phase diagram of the $\mathbb{Z}_{2}$ gauge–Higgs model in 3D, showing the confinement transition line (red), the Higgs transition line (black), the self-dual multicritical point (MCP) where they meet, and a segment of first order coexistence extending from the MCP (double line). The model’s partition function is shown in Eq. 1, with $x,y$ defined in Eq. 2. The partition function is expressed in terms of classical membranes in Eq. 3, where $-\log x$ is the bare membrane tension and $-\log y$ is the bare line tension of membrane boundaries.
@@
In the 3D $\mathbb{Z}_{2}$ gauge-Higgs model, the Higgs and confinement lines (shown in Fig. 1) are completely equivalent, by the exact duality property of the model [1]. In our convention, the transition line we discuss is the confinement transition. As can be seen in Fig. 1 we can cross this line at various values of the Higgs coupling, which is related to $y$ in the figure. As we move along the confinement phase transition line by increasing $y$, the typical size $\ell$ of the “holes” in the membrane increases, diverging at $y_{\text{MCP}}$, the self-dual critical point where Higgs and confinement lines meet [14, 20]. We study the confinement (Ising) transition at various values of $y$, up to and including $y_{\text{MCP}}$.
@@ II.1 Definition of the model
There are two dual ways to formulate the membrane mapping. The equivalence of these dual representations is one way to see the equivalence of the Higgs and confinement phase transition lines in (Fig. 1). In a quantum language, these dual formulations correspond to formulations of the path integral either in the “electric” or the “magnetic” basis, as discussed below. We choose the “electric” representation, in which the membrane ensemble is [^3] The electric representation is obtained by a strong coupling (“high temperature”) expansion of Eq. 1. The dual (i.e. “magnetic”) representation is an even more straightforward rewriting of Eq. 1: the membranes live on the dual cubic lattice, and a plaquette of the dual cubic lattice is occupied if ${\tau_{i}\sigma_{ij}\tau_{j}=-1}$ for the corresponding link $\left\langle ij\right\rangle$ of the original lattice. See e.g. App. A of [14] for more detail.
@@
Eqs. 1 and 3 can be thought of as 3D classical partition functions in their own right. However, if we choose to think of $Z$ as an imaginary time path integral for a 2+1D quantum $\mathbb{Z}_{2}$ gauge theory, then the membrane ensemble (3) is the formulation of this path integral in the basis of $\mathbb{Z}_{2}$ electric flux lines which live in the 2D spatial plane and which terminate on electric charges. The membranes we are discussing are worldsurfaces of these electric flux lines (see e.g. Fig. 6 in Ref. [14]). The boundaries of the membranes, which make up the “holes”, are worldlines of the electric charges. Duality is the existence of a completely equivalent picture in the magnetic basis, for $\mathbb{Z}_{2}$ magnetic flux lines and charges.
@@ II.3 Finite and infinite holes
Heuristically, when $\ell(y)$ is finite (${y<y_{\text{MCP}}}$), we might imagine that holes disappear under RG. That is, after coarse-graining to scales ${\gg\ell(y)}$ we effectively recover a theory of closed membranes, and the system flows to the same Ising fixed point that governs the phase transition on the boundary ${y=0}$. Below we will show that this can be made concrete (giving a constructive mapping to Ising) also for ${y>0}$.
@@ III.2 Membrane-patching algorithm
We have proposed an explicit, easy to implement, patching procedure. Other legitimate procedures, leading to slightly different patched surfaces, are also possible, but such changes would not affect the universal behavior of the correlation function at distances $r\gg\ell$ (Sec. III.4, Sec. V), or the analysis of the phase transition. Having defined the patches we can define the repaired configuration $\overline{\mathcal{M}}$. If we call the set of patching plaquettes ${\cal P}$, the new set ¯$\overline{\mathcal{M}}={\cal M}\cup{\cal P}-{\cal M}\cap{\cal P}$ defines strictly closed surfaces and allows us to construct line operators as described in Sec. III.1.
@@ III.3 Ising observables
As discussed around Eq. 4, an ensemble of closed membranes can be split into sectors. Our patching algorithm works in all sectors. It is also perfectly meaningful to define correlation functions of the $\widetilde{V}$ operators conditioned on a given sector (as defined more precisely below): correlators for different sectors agree when $r\ll L/2$, so that restricting to a sector does not affect the location of the phase transition, but may differ when $r$ is of order $L$.
One reason it is often useful to restrict to a sector is for the numerical analysis of the confinement phase transition. The analysis of the transition through the asymptotics of the correlators is subject to large statistical errors; more effective procedures are based on the analysis of Binder-like parameters [30]. These can be defined straighforwardly in the first sector (${\sigma_{x}=\sigma_{y}=\sigma_{z}=+}$).
In this sector the dual Ising spin $S_{r}$ has periodic boundary conditions, and so is well-defined up to a global sign flip: this means that, after patching, the correlation function $\left\langle S_{r}S_{r^{\prime}}\right\rangle$ may be straightforwardly defined. When we write expressions such as $\left\langle S_{r}S_{r^{\prime}}\right\rangle$, we refer to an average conditioned on the ${\sigma_{x}=\sigma_{y}=\sigma_{z}=+}$ sector.
More formally, this may be expressed as
@@
where $P$ is any path connecting $r$ to $r^{\prime}$, and $\chi_{+}(\mathcal{M})$ is an indicator function that takes the value $\chi_{+}=1$ if $\mathcal{M}$ is a patchable configuration in the +++++++ + + sector, and the value $\chi_{+}=0$ otherwise.
@@
III.4 Diagnosing the deconfined phase;
Robustness of patching procedure
Since in Sec. IV below we will focus on critical scaling near the confinement transition line (the red line in Fig. 1), we emphasize in this section that the patching procedure gives a way to detect or verify deconfinement more generally (even far away from the confinement transition line).
First we note that the conditions ($i$) and ($ii$) below serve as (easily checkable) sufficient conditions for a point of the phase diagram to lie within the deconfined phase. Subsequently we will argue that, for the present patching algorithm and model, they are also necessary conditions (meaning that patching can be used as a diagnostic of deconfinement throughout the entire deconfined phase):
$\bullet$ Condition ($i$): The patching procedure succeeds: i.e. the probability of a configuration $\mathcal{M}$ being patchable tends to 1 at large system size. Equivalently, the holes are non-percolating, as discussed below.
This condition ensures that the line operators $\widetilde{V}$ that we define through patching are quasilocal (as discussed in more detail in Sec. V). Given this, these operators may then be used to detect deconfinement.
In principle, one way to detect deconfinement would be by checking for the spontaneous breaking of one-form symmetry using a line operator that wrapped around the system in some direction. An alternative (and more practical [^7] In the quantum language, a signature of the deconfined phase is that a line operator $\hat{\widetilde{V}}$, which winds around the system in some direction, acts nontrivially on the ground state space. As a result $\langle\widetilde{V}\rangle=0$ for the thermal expectation value of a winding string operator, so long as the thermal expectation value properly sums over all topological sectors (Sec. II.2). However, because of the order–$L$ free energy barrier that separates different topological sectors (in the 3D classical language), it will often be much easier for Monte Carlo to sample a single sector than to sample all sectors. Therefore observables that can be calculated in a single sector are more useful in practise. This is not a serious restriction since it is possible to detect the thermodynamic phase using observables within a topological sector.) diagnostic is the expectation value of an open string operator:
$\bullet$ Condition ($ii$): The expectation value of the string operator ${\lim_{L\to\infty}\langle\widetilde{V}_{P}\rangle}$, for an open path $P$ of length $R$, tends to zero at large $R$.
In terms of the dual Ising degrees of freedom, this means that the dual Ising order parameter is disordered. Numerically, a convenient way to detect this is via the vanishing at large $L$ of the squared magnetization $\langle M^{2}\rangle$ (defined in the previous section).
The two conditions above show that patching gives a simple diagnostic for the deconfined phase: first we confirm that holes do not percolate, and then we check that $\langle M^{2}\rangle$ vanishes at large $L$.
If, in the second step, we find that $\langle M^{2}\rangle$ does not vanish, this tells us that we are not in the deconfined phase. This is what happens if we are to the left of the confinement line in Fig. 1.
The remaining question is about the nature of the phase transitions where the holes start to percolate, i.e. where condition ($i$) ceases to hold. A priori, there are two possibilities. First, the percolation transition of the holes could happen at the thermodynamic phase boundary of the deconfined phase (the Higgs transition line in Fig. 1, together with the multicritical point). Alternately, the percolation transition line could be separated from the thermodynamic phase transition line, occurring at a smaller value of $y$ that had no thermodynamic significance.
For the present model, the former scenario holds. In other words, the Higgs transition line is the percolation phase transition of the holes. This was already anticipated in Ref. [20]; Ref. [14] gave numerical evidence that the hole percolation transition coincides with the thermodynamic transition along the full extent of the Higgs line (plus multicritical point), and sketched an analytical explanation for why the two transitions coincide which we develop a little further in App. B.
As a result, the present patching approach successfully diagnoses the entirety of the deconfined phase. Further, we argue in App. B that this property is robust to slight changes of the model or the patching scheme.
On the other hand, a sufficiently severe change to the way clusters are defined in the patching scheme (a sufficiently “bad choice” of patching algorithm) could cause percolation to occur in the interior of the deconfined phase, as we discuss in App. B. In that scenario, the “bad” patching algorithm would still be able to successfully diagnose deconfinement for small enough $y$, but for larger $y$ it would not give us any information about which phase the model was in.
@@ IV.1 Dual Ising correlators
A striking demonstration of Ising universality is given by the power-law scaling of the dual Ising correlator as a function of separation $z$:
@@
The correlator is shown in Panel 1 of Fig. 5, as a function of distance $z$ in a system of fixed size $L=64$. Data for the values ${y=0,0.1,0.2}$, which lie on the confinement line and are expected to be in the Ising universality class, are consistent with the Ising power law, where $\eta_{\mathrm{Ising}}=0.036298(2)$ [31], which is shown with a dotted line. [^8] The power law applies for $1\ll r\ll L$. It is also possible to perform a scaling collapse for $r/L$ of order 1 (not shown). (Since the Ising anomalous dimension $\eta_{\text{Ising}}$ is small, this line is close to $1/z$.) The distinct curves are also consistent with each other within error bars: although in general the nonuniversal normalization of $C(z)$ will depend on $y$, this dependence appears to be very weak. The Panel also shows the correlator computed at the MCP, where its RG interpretation is different and will be discussed in Sec. V.2.
@@ IV.2 Other sectors
At the Ising critical point, and in the limit $L\rightarrow\infty$, these numbers are universal and related by duality to ratios of Ising partition functions with different boundary conditions,
@@
In our Monte Carlo dynamics, the configuration can make a transition between different sectors by passing through a non-patchable configuration with a percolating hole. In principle this allows equilibration over all sectors. In practice this is possible if $y$ is not too small and $L$ is not too large. When $y<y_{\text{MCP}}$ such transitions are exponentially suppressed in the limit $L\rightarrow\infty$ simply because $P_{\text{perc}}(L)$ is exponentially suppressed. For this reason we have not tried to accurately compute the sector probabilities for the Ising critical point. This could be done, using the present gauge theory, by parallel tempering, exchanging configurations at different values of $y$ [32].
@@ V.1 Case where $\ell$ is finite
An alternative way to quantify locality in the dual language would be to reexpress the partition function in terms of $S_{r}$.121212 Separating $Z$ into topological sectors (see previous sections) we write e.g. $Z_{+++}=\sum_{\mathcal{M}\in(+++)}e^{-\mathcal{H}[\mathcal{M}]}$ (for the +++++++ + + sector; other sectors are similar) as $Z_{+++}=\frac{1}{2}\sum_{\{S_{r}\}}\sum_{\mathcal{M}\in(+++)}e^{-\mathcal{H}[\mathcal{M}]}\chi(S,\mathcal{M})$, where $\chi(S,\mathcal{M})=1$ if $S$ is one of the two spin configurations consistent with the patched configuration $\overline{\mathcal{M}}$ and zero otherwise. Then the effective spin Hamiltonian is given by $e^{-\mathcal{H}^{\prime}[S]}=\sum_{\mathcal{M}}e^{-\mathcal{H}[\mathcal{M}]}\chi(S,\mathcal{M})$. Because of patching the effective interactions would be nonlocal, but could be approximated by interactions of finite range $\mathcal{O}(\ell)$ at the cost of incurring only a small error in the free energy density.